OpenStudy (kenww):

I have no clue how to do this one. Find the specified vector or scalar. u = -4i + 1j and v = 4i + 1j; Find ‖u+v‖.

OpenStudy (kenww):

yes that is all the info given

OpenStudy (amorfide):

\[u=-4i+1j\] \[v=4i+1j\] \[||u+v||=\sqrt{(u_{1}^{2}+v_{1}^{2})+(u_{2}^{2}+v_{2}^{2})}\] so do u+v u+v=(-4i+4i) + ( 1j+1j)=0i + 2j = 2j so the magnitude of this vector becomes \[||u+v||=\sqrt{(0)^{2} + (2)^{2}}\]

OpenStudy (kenww):

so that is the specified vector?

OpenStudy (amorfide):

the specified vector or scalar, is ||u+v||, so yes

OpenStudy (amorfide):

you can finish the rest

OpenStudy (kenww):

right thanks